Shift report 20030129 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 29 2003 - 00:12:29 EST

  • Next message: Apache: "Shift report 20030129 00:00-08:00"
    ------------------------------------------------------------
    Date: Wed Jan 29 00:12:29 2003, Shift: 16:00-24:00
    Supervisor: PS, with: 
    Log book#13 pages 133 - 134
    
    Shift Summary:
    ------------------------------------------------------------
    
    1. Beam from 17:00 - 20:00.
       Later tried twice to restore the beam but not succeeded.
    
    2. Data runs: 7685, 7686, 7687 and 7688 - about 300k #6,
    but for runs 7687-7688 the backgroud was very high. 
    7689 - pedestal run,
    7690 - TPC pedestal run.
    
    3. No detector problems.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed Jan 29 2003 - 00:13:00 EST